OpenStudy (anonymous):

The numbers 1 to 40 are marked on ping-pong balls and put into a barrel. If a ball is drawn at random, find: (i) P(multiple of 4) (ii) P(multiple of 7) (iii) P(multiple of 28) (iv) P(multiple of 4 or 7). (b) Are the events ‘drawing a multiple of 4’ and ‘drawing a multiple of 7’ independent? Explain. I just need help with the last question - B - but I thought the other information would be useful.

OpenStudy (anonymous):

there are 40 balls in all.. since there are 10 numbers between 1 and 40 inclusive that are multiples of 4, P(mult4) = 10/40 = 1/4 since there are 5 numbers between 1 and 40 inclusive that are multiplse of 7, P(mult7) = 7/40 since there is only 1 number between 1 and 40.. blah. blah blah., P(mult28) = 1/40 since there are 10 #s mult of 4 and 7 #s mult 7 and since 28 is both a member of mult4 and mult7, you only need to count that once... so P(mult4 or mult7) = 16/40 = 2/5